Status Report

NASA Space Station On-Orbit Status 5 May 2004

By SpaceRef Editor
May 5, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 5 May 2004

All ISS systems continue to function nominally, except those noted previously or below.

Both CDR Gennady Padalka and FE/SO Michael Fincke in turn took their first periodic on-orbit hearing assessment (O-OHA) test, a NASA environmental health systems examination to assess the efficacy of acoustic countermeasures.   [The O-OHA audiogram test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special “EarQ” software on the MEC (medical equipment computer). The baseline test is required not later than about Flight Day 14 for each new Expedition and is then performed once per month.]

As is standard practice after Russian dockings, CDR Padalka took two photos of the internal FGB (funktsionalnyi-grusovoi blok) module nadir docking assembly (StA) used for the Soyuz TMA-4/8S linkup. These images will be used to refine current understanding of docking conditions.   [The ASP-B passive docking assembly of the FGB (-y axis) is used for docking transport and resupply vehicles. The objective was to take photo imagery of the scratch or scuff mark left by the head of the docking probe on the internal surface of the StA’s receiving cone. As other cosmonauts before him, Gennady used the Kodak 760 digital still camera to take two pictures with the hatch closed down and downlinked them later (5:35am EDT) via OCA.]

Science Officer Fincke recharged both batteries of the MedOps defibrillator equipment to full capacity.   [Each battery was charged for ~3.5 hrs, and its open-circuit voltage was tested with the “Pilobolus” volt/amp scopemeter, then removed and stowed again. Nominally, the Defib has a battery installed at all times, but with this particular unit the PDIM (power data interface module) is not functioning properly and would overcharge the batteries if left inside. They have to be charged every 60 days along with the Defib checkout.]

Systems monitoring of the Soyuz TMA-4 continues. Twice today, the CDR took routine spacecraft parameter readings (Form 03) in the vehicle for calldown to TsUP/Moscow.

After switching filter bed 1 of the SM’s harmful impurities removal unit (BMP) yesterday to Regeneration cycle, Gennady today terminated the bake-out, moding the channel back to Purify. Later, he initiated regeneration on filter channel 2. [The regeneration of the air purifier filter beds is repeated every 20 days. Each bakeout to space vacuum takes about 24 hours.]

Mike Fincke completed the seventh imaging session of the BCAT-3 (Binary Colloidal Alloy Test-3) experiment, setting up the SGSM (slow growth sample module) on the MWA (maintenance work area), photographing the samples (today #3 & #6) sequentially with the Kodak 760 digital still camera and then stowing them, with the data stored on a PCMCIA 1GB micro drive. Afterwards, the BCAT MWA hardware was disassembled and stowed.

Fincke terminated the maintenance charging/discharging cycle on EMU/spacesuit batteries in the Airlock’s BSA (battery stowage assembly), initiated two days ago on batteries #2045 & #2046.   [The automatically controlled charging took about 24 hrs and was followed by discharge. Helmet light and PGT (pistol grip tool) batteries were not charged at this time and had to be removed from the BSA beforehand. After the procedure, Mike stowed the batteries in the Airlock, with their 50-day maintenance clock reset.]

Mike also conducted routine maintenance on CSA-CPs (compound specific analyzer-combustion products) #1003, #1009 and #1010. Later, he used all three units for an air data take.

Gennady Padalka continued the latest round of periodic preventive maintenance of Russian segment (RS) ventilation systems, today cleaning the VD1 and VD2 air ducts in the “Pirs” DC-1 docking compartment. (Last time done: 3/30).

Mike Fincke powered up the MSG (Microgravity Science Glovebox) rack for an untended ground-commanded facility checkout of thermal sensor data without MTL (moderate temperature loop) flow.   [The activity, required for MSG ops recertification, involved ESEM (exchangeable standard electronic module) activations, delta pressure & humidity sensor readings, and downlinking (with voice) of facility resource utilization data.]

The two crewmembers conducted a 30-min. familiarization checkout with the CMRS (crew medical restraint system), unstowed yesterday from the CHeCS (crew health care systems) rack.   [The board-like CMRS allows strapping down a patient on the board with a harness for medical attention by the CMO (crew medical officer) who is also provided with restraints around the device. CMRS can be secured to the ISS structure within two minutes to provide a patient restraint surface for performing emergency medical procedures, such as during ACLS (advanced cardiac life support). It can also be used to transport a patient between the station and the Orbiter middeck. It isolates the crew and equipment electrically during defibrillations and pacing electrical discharges, accommodates the patient in the supine zero-G positions, provides cervical spine stabilization and, for a three-person crew, can also restrain two CMOs (crew medical officers) during their delivery of medical care.]

On the HRF (Human Research Facility) rack, the SO reconfigured laptop settings and installed a number of audio drivers (from Inc. 8) to support upcoming ADUM OPE (advanced ultrasound/on-board proficiency enhancer) operations.

Padalka performed the periodic replenishing of the Elektron’s water supply for electrolysis, filling the KOV thermal loops’ EDV container with purified water from the BKO multifiltration/purification column unit.   [The procedure is specially designed to prevent air bubble formation that may cause micropump impeller cavitation and Elektron shutdown.]

Mike Fincke conducted the daily leak check of the Lab window’s inter-pane volume, using the “Aeolus” scopemeter with pressure probe. [Past readings, dating back to March, have found a steady leak rate of ~27 Torr (0.52 psi) per day from the cabin into the interstitial volume (i.e., not overboard).]

In the SM, Gennady conducted the routine technical maintenance (including ASU toilet system replacements) of the SOZh environment control & life support system and performed the preparation of the daily equipment stowage “delta” file for updating the IMS (inventory management system) database, while Mike attended to the regular routine status checkup of autonomous Lab payloads.

The crew completed their daily physical exercise program on TVIS treadmill, RED exerciser, CEVIS ergometer and VELO cycle with load trainer. Fincke also did the monthly maintenance of the CEVIS, mainly involving an examination of the wire rope isolators for damage.

Both crewmembers again had an hour each for general station familiarization and acclimatization, as is standard daily rule for the first two weeks after starting station residency.

Working off the Russian task list, Padalka prepared the EGE-2 laptop and the Halley payload equipment for the Neurocog VC5 “Cervantes” Neurocog experiment scheduled for tomorrow.   [By researching the integration of visual, vestibular, and proprioceptive cues in the perception of body position in space, Neurocog studies how humans perceive space, what role the sensory information of sight, balance, motion and position plays in this, and how human perception is affected by weightlessness.]

After attitude control handover from the U.S. CMG Momentum Manager system to Russian MCS (motion control system) thrusters at 10:40am EDT, the ISS at 10:54am performed a 90-degree turn from LVLH XVV (local vertical local horizontal/x-axis in velocity vector) to LVLH YVV (y-axis in velocity vector), i.e., flying “sideways”. Control authority was returned to U.S. CMGs at 11:25am directly (i.e., without intermediate desat attitude hold).   [According to the new CMG-protective strategy, the station will remain in this attitude until 5/18.]

A software transition of the two EXT MDMs (exterior multiplexer/demultiplexers, computers) on the S0 truss segment to the new load, version EXT R4, is in pre-approval planning. The software transition is tentatively scheduled for next week (5/11-14).

Onboard resources continue to be monitored carefully. All supplies (water, N2, O2, food, propellant, TVIS bearings, RED canisters) are tracking to predictions.   [Current ISS propellant quantity is 3725 kg, after last week’s use of 59 kg for 7S undocking and attitude maneuvers.]

Starting at 4:00am EDT tomorrow morning and running for seven hours, MCC-H and its support group in Moscow (HSG) will conduct another BCC (backup control center) dry run in test mode, with no involvement of the ISS crew or vehicle.

On Monday (5/3), battery #6 of the SM went off line when its Charge/Discharge Unit (ZRU) was deactivated due to a trip of its Limit Discharge Sensor (DPR).   [SM power is being supplied by the remaining seven 800A batteries and SNTs (voltage & current stabilizers/transformers) until the DPR issue is resolved.]

TsUP/Moscow continues long-term testing to troubleshoot the Russian ASN satellite navigation system aboard the SM.   [Proper functioning of the ASN will be critically important for the docking of the European ATV (automated transport vehicle).]

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Muglad Basin Fans, Sudan (good pass over the west end of the basin, so that views from nadir and panoramas looking left as far as the Nile, should have captured this critically interesting zone. The broad fans are clustered next to each other and are probably an analog of the deeply buried landscapes which now host oil directly beneath), Ubangi River, Congo (Dynamic event. Congo basin was unusually clear: looking right down the Ubangi River), Congo River mouth, Congo (Dynamic event. Looking a touch left and nadir for the mouth of the Congo, and right for clear views up into the usually cloudy basin), Tucson, Arizona (nadir pass), and Lake Poopo, Bolivia (Lake Poopo is not dry at the moment. Its level varies directly with El Nino cycles. The almost full lake can be detected, left of track, between Lake Titicaca and the big white salt lake, Salar Uyuni).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status  (as of today, 12:45pm EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Regeneration mode. RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20). SFOG slot#2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 24.9; ppO2 (mmHg) — 155.6; ppCO2 (mmHg) — 3.2;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.3.
  • FGB Cabin:  Pressure (mmHg) — 759; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 747.71; temperature (deg C) — 22.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 749.97; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 750.17; temperature (deg C) — 22.7; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed position (blind, dual-angle mode with drag reduction, non solar-tracking).
  • SM batteries:  Battery #6 is off line; battery #1 is in Cycle mode (RTs); all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #1 is off-line (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3725 (8212 lb) as of 5/5/04;  [SM(693) + FGB(2373) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2 RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:28am EDT [= epoch]):

  • Mean altitude — 361.1 km
  • Apogee — 368.7 km
  • Perigee — 353.6 km
  • Period — 91.77 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0011145
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 31161

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.